OpenStudy (anonymous):

The equation of a line is x + 5y = 12. What is the y-intercept of the line? (Points : 1)

OpenStudy (anonymous):

Option A: -12 Option B: 12 Option C: 5/12 Option D: 12/5

OpenStudy (kkutie7):

solve for y first.

OpenStudy (kkutie7):

\[x+5y=12\] \[5y=12-x\] \[y=\frac{12}{5}-\frac{x}{5}\] move into y=mx+b form where b is your y-int \[y=\frac{-x}{5}+\frac{12}{5} \]

OpenStudy (kkutie7):

\[b=\frac{12}{5}\]